OpenStudy (anonymous):

I need help with this question.... Name the algebraic property demonstrated in the example below: (2 + x) + y = 2 + (x + y) (1 point) Commutative Property of Addition Associative Property of Addition Distributive Property Transitive Property

OpenStudy (calculusxy):

One would usually associate this with commutative property, HOWEVER, since this has parenthesis (which does make a huge difference in some cases) it would be associative property.

OpenStudy (anonymous):

thank you

OpenStudy (calculusxy):

You're welcome :)

OpenStudy (anonymous):

could you help with another question?

OpenStudy (calculusxy):

Can try...

OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below: x • y • z = y • x • z (1 point) Distributive Property Transitive Property Associative Property of Multiplication Commutative Property of Multiplication

OpenStudy (calculusxy):

This one would be commutative because the terms are just switched around but it mean the same thing.

OpenStudy (anonymous):

but how

OpenStudy (calculusxy):

Just another thing to throw off there: if this had a parenthesis then it would be the associative property of multiplication.

OpenStudy (anonymous):

ohhhh ok thx :)

OpenStudy (calculusxy):

"The word "commutative" comes from "commute" or "move around", so the Commutative Property is the one that refers to moving stuff around. For addition, the rule is "a + b = b + a"; in numbers, this means 2 + 3 = 3 + 2. For multiplication, the rule is "ab = ba"; in numbers, this means 2×3 = 3×2." -- purplemath.com

OpenStudy (calculusxy):

No problem :)